# Temporal Inductive Path Neural Network for Temporal Knowledge Graph Reasoning

### 08 Dec 2023 v2

@@ 5.1.2. Evaluation Metrics
In contrast to the traditional filtered setting used in previous works (Bordes et al., 2013)(Jin et al., 2020)(Zhu et al., 2021a), where all valid quadruples appearing in the training, validation, or test sets are removed from the ranking list of corrupted facts, we consider that this setting is not suitable for TKG reasoning tasks. Instead, we opt for a more reasonable evaluation, namely time-aware filtered setting, where only the facts occurring at the same time as the query are filtered from the ranking list of corrupted facts, which is aligned with recent works (Sun et al., 2021b)(Han et al., 2021)(Dong et al., 2023).
5.1.3. Compared Methods
@@
We conduct a comprehensive comparison of our proposed model with three categories of baselines: (1) KG Reasoning Models. It consists of traditional KG reasoning models that do not consider timestamps, including DistMult(Yang et al., 2014), ComplEx (Trouillon et al., 2016b), ConvE (Dettmers et al., 2017), and RotatE (Sun et al., 2019). (2) Interpolated TKG Reasoning models. We also compare our proposed model with four interpolated TKG reasoning methods, namely TTransE (Leblay and Chekol, 2018), TA-DistMult (Garcia-Duran et al., 2018), DE-SimplE (Goel et al., 2020), and TNTComplEx (Lacroix et al., 2020). (3) Extrapolated TKG Reasoning models. We include state-of-the-art extrapolated TKG reasoning methods that infers future missing facts, including TANGO-Tucker (Han et al., 2021), TANGO-DistMult (Han et al., 2021), CyGNet (Zhu et al., 2021a), RE-NET (Jin et al., 2020), RE-GCN (Li et al., 2021), TITer (Sun et al., 2021a), xERTE (Han et al., 2020a), CEN (Li et al., 2022b), GHT (Sun et al., 2022), and DaeMon (Dong et al., 2023).
5.1.4. Implementation Details
@@
For history temporal graph construction, we perform a grid search on the history length $m$ and present overview results with the lengths 25, 15, 10, 8, corresponding to the datasets ICEWS18, GDELT, WIKI, and YAGO in Table 3, which is described in detail presented in Figure 7. The embedding dimension $d$ is set to 64 for temporal path representation. For query-aware temporal path processing, we set the number of temporal path aggregation layers $\omega$ to 6 for ICEWS18 and GDELT datasets, and 4 for WIKI and YAGO datasets. We conduct layer normalization and shortcut on the aggregation layers. The activation function $relu$ is adopted for the aggregation of the temporal path. Similar to (Zhu et al., 2021b), we use different temporal edge representations in different aggregation layers, that is the learnable parameter in $\bm{\Psi}$ and $\bm{\Upsilon}$ is independent of each aggregation iteration (Sadeghian et al., 2019) (Yang et al., 2017). To facilitate parameter learning, we have set the number of negative sampling to 64 and the hyperparameter $\alpha$ in the regularization term to 1. We use Adam (Kingma and Ba, 2014) for parameter learning, with a learning rate of ${1e-4}$ for YAGO and ${5e-4}$ for others. The maximum epoch of training has been set to 20. We also design an experiment in an inductive setting, which predicts facts with unseen entities in the training set, to demonstrate the inductive ability of our proposed model. All experiments were conducted with EPYC 7742 CPU and 8 TESLA A100 GPUs.
5.2. Experimental Results (RQ1)
@@
Table 3. Overall Performance Comparison of Different Methods. Evaluation metrics are time-aware filtered MRR and Hits@{1,3,10}. All results are multiplied by 100. The best results are highlighted in bold. And the second-best results are highlighted in underline. (Higher values indicate better performance.)
@@ 5.3. Ablation Study (RQ2)
5.3.1. Effectiveness of Temporal Relation Encoding
@@
Figure 4. Ablation Results on Temporal Encoding.
@@
Table 4. Impact of the Independence Setting of Parameters in Temporal Encoder.
@@
5.3.2. Variants of Temporal Edge Merging Method
The temporal edge merging operation $\textsc{Tmsg}(\cdot)$ utilizes the message passing mechanism to iteratively extend the path length while merging the query-relevant path information. As mentioned in Equation 6, we employ the scaling operator from DistMult (Yang et al., 2014) for computing the current temporal path representation with the features of temporal edges through element-wise multiplication. Additionally, we have also conducted experiments using the translation operator from TransE (Bordes et al., 2013) (i.e., element-wise summation) and the rotation operator from RotatE (Sun et al., 2019) to provide further evidence supporting the effectiveness of the temporal edge merging operation.
@@ 5.3.2. Variants of Temporal Edge Merging Method
Figure 5. Ablation Results on Temporal Edge Merging Method.
5.3.3. Variants of Path Aggregation Method
The path aggregation operation $\textsc{Agg}(\cdot)$ is performed after merging the temporal edges at each step, and the aggregated feature serves as the temporal path representation from the query subject entity to the object entity candidates for future fact reasoning. As mentioned in Section 4.4.2, we employ the principal neighborhood aggregation (PNA) (Corso et al., 2020) as the aggregator to aggregate the propagated messages. PNA aggregator considers mean, maximum, minimum, and standard deviation as aggregation features to obtain comprehensive neighbor information. To validate its effectiveness, we compare it with several individual aggregation operations: SUM, MEAN, and MAX.
@@ 5.3.3. Variants of Path Aggregation Method
Figure 6. Ablation Results on Path Aggregation Method.
5.4. Parameter Study (RQ3)
@@
5.4.1. Sampling History Length
In order to comprehensively capture the interconnection patterns between entities in the historical subgraphs, we constructed a history temporal graph to learn query-aware temporal paths between entities, enabling the prediction of future missing facts. Unlike previous methods based on subgraph sequences (Li et al., 2021) (Li et al., 2022b), we fuse multiple historical subgraphs for reasoning. However, it may lead to the inclusion of redundant edge information from different timestamps when the historical length is excessively long, resulting in unnecessary space consumption in the learning process. Moreover, facts that are far from the prediction target timestamp have minimal contribution to the prediction, even with the addition of relative distance-aware temporal encoding in the temporal relation encoder. Conversely, if the history length is too short, it can hinder the model’s ability to capture cross-time features between entities, thereby reducing the modeling performance for temporal patterns in the history temporal graph. Hence, we conducted a discussion on the sampling history length to find a balance. The experimental results of the history length study are shown in Figure 7.
@@ 5.4.1. Sampling History Length
Figure 7. The Performance of Different History Length Settings.
Table 5. Density and Interval of Datasets ($|\mathcal{V}_{avg}|$ is the average number of entities involved in the subgraph of each timestamp.).
@@
5.4.2. Number of Temporal Path Aggregation Layers
The temporal path aggregation layer is a fundamental computational unit in TiPNN. It operates on the constructed history temporal graph using the temporal edge merging and path aggregation operation to learn query-aware temporal path representations for future fact inference. The number of layers in the temporal path aggregation layer directly impacts the number of hops in the temporal message passing on the history temporal graph, which corresponds to the logical maximum distance of temporal paths. Therefore, setting the number of layers faces a similar challenge as determining the history length in Section  5.4.1.
@@ 5.4.2. Number of Temporal Path Aggregation Layers
Figure 8. The Performance of Different Numbers of Aggregation Layers.
@@
5.5. Comparison on Prediction Time (RQ4)
To further analyze the efficiency of TiPNN, we compared the inference time of TiPNN with DaeMon (Dong et al., 2023) on the TKG reasoning task. For a fair comparison, we use the test sets of four datasets and align the model parameters under the same setting and environment. The runtime comparison between the two models is illustrated in Figure 9, where we present the ratio of their inference times in terms of multiples of the unit time. The comparative experiments demonstrate that TiPNN achieves significant reductions in runtime compared to DaeMon, with time savings of approximately 80%, 69%, 67%, and 73% on the ICEWS18, GDELT, WIKI, and YAGO datasets, respectively.
Figure 9. Runtime Comparison (The runtime is proportionally represented as multiples of the unit time.).
@@ 5.5. Comparison on Prediction Time (RQ4)
5.6. Inductive Setting (RQ5)
To validate the model’s inductive reasoning ability, we also design experiments for TiPNN in an inductive setting.
The inductive setting is a common scenario in knowledge graphs, where during the training process, the model can only access a subset of entities, and during testing, it needs to reason about unseen entities. This setting aims to simulate real-world situations where predictions are required for new facts based on existing knowledge. In the context of the temporal knowledge graph, the inductive setting requires the model to perform reasoning across time, meaning that the learned representations from the historical subgraphs should generalize to future timestamps for predicting missing facts. This demands the model to possess the ability to generalize to unseen entities, leveraging the learned temporal connectivity patterns from the historical subgraphs.
@@ 5.6. Inductive Setting (RQ5)
In the context of TKGs, there are few existing datasets suitable for inductive validation. Therefore, we follow the rules commonly used in KG for inductive settings and construct an inductive dataset specifically tailored for TKGs (Teru et al., 2020). Finally, we conduct experiments on our proposed model using the inductive dataset to evaluate its inference performance under the inductive setting.
5.6.1. Inductive Datasets
@@
Table 6. Statistics of Inductive Datasets ($\mathcal{E}_{train}$, $\mathcal{E}_{valid}$, $\mathcal{E}_{test}$ are the numbers of facts in training, validation, and test sets.).
@@
5.6.2. Results of Inductive Experiment
@@
Table 7. Inductive Performance of TiPNN ($\rightarrow$ points from training set to test set corresponding to the inductive setting, and the row results without arrow marked is the corresponding transductive setting as the baseline. ‘Bias’ denotes represents the difference in comparison.).
@@
5.7. Reasoning Evidence (RQ6)
Since we have integrated historical information into the constructed history temporal graph and perform inference on future facts by modeling temporal paths, we can use the path in the history temporal graph to provide evidence for the inference process and visualize the reasoning basis. An intuitive idea is that the model should provide important reasoning paths from the history temporal graph that contribute significantly to the inference for the corresponding response. Although the temporal path representation we obtain is a comprehensive representation logically aggregated from multiple paths in the history temporal graph, we can still estimate the importance of each individual path following the local interpretation method (Baehrens et al., 2010) (Zeiler and Fergus, 2014). As described in Section 4.5.1, we convert the learned temporal path representation into corresponding scores, and thus, we can estimate the importance of paths through backtracking.
Table 8. Reasoning Evidence of Responses to Selected Queries from ICEWS18. The queries, responses, and the corresponding top-2 reasoning evidence with importance (higher values indicate higher importance) are provided. The attached superscript ${}^{-1}$ denotes the inverse relations. The timestamp in the quadruple is converted into mm/dd/yyyy format for representation.
@@ 5.7. Reasoning Evidence (RQ6)
0.784 ¡Shinzo Abe, Express intent to meet or negotiate, China, 10/22/2018¿.
0.469 ¡Shinzo Abe, Express intent to meet or negotiate, Li Keqiang, 10/21/2018$\rightarrow$
@@
¡Shinzo Abe, Make a visit, China, 10/22/2018¿.
@@
1.613 ¡European Union, Engage in diplomatic cooperation, United Kingdom, 10/14/2018¿.
1.044 ¡European Union, Express intent to engage in diplomatic cooperation (such as policy support),
United Kingdom, 10/22/2018¿.
@@
0.376 ¡Citizen (Thailand), Accuse, Thailand, 10/03/2018$\rightarrow$
¡Thailand, Return, release person(s), Citizen (Thailand), 10/10/2018$\rightarrow$
@@
0.757 ¡Health Ministry (India), Make an appeal or request, Citizen (India), 10/26/2018¿.
0.367 ¡Health Ministry (India), Make statement, Government (India), 10/19/2018$\rightarrow$
@@
Drawing inspiration from path interpretation (Zhu et al., 2022), we define the importance of a path as the weights assigned to it during the iteration process. We achieve this by computing the partial derivative of the temporal path scores with respect to the path’s weights. Specifically, for a reasoning response $(s,r,o,t+1)$ of a missing object query $(s,r,?,t+1)$, we consider the top-k paths as the basis for inference, which is defined as shown in Equation 16.
(16) $\text{P}_{1},\text{P}_{2},\cdots,\text{P}_{k}=\underset{\text{P}\in\mathcal{P}_{s\rightarrow o}^{\prec t+1}}{\text{Top-}k}\frac{\partial\ p(s,r,o)}{\partial\ P}$
@@
6. Conclusion
In this work, we have introduced TiPNN, an innovative query-aware temporal path reasoning model for TKGs reasoning tasks, addressing the challenge of predicting future missing temporal facts under the temporal extrapolated setting.

### 25 Jan 2024 v3

@@ 4.1.2 Evaluation Metrics
In contrast to the traditional filtered setting used in previous works bordes2013translating ; jin2020Renet ; zhu2021learning , where all valid quadruples appearing in the training, validation, or test sets are removed from the ranking list of corrupted facts, we consider that this setting is not suitable for TKG reasoning tasks. Instead, we opt for a more reasonable evaluation, namely time-aware filtered setting, where only the facts occurring at the same time as the query are filtered from the ranking list of corrupted facts, which is aligned with recent works 2021TimeTraveler ; han2021learning ; dong2023adaptive .
4.1.3 Compared Methods
@@
We conduct a comprehensive comparison of our proposed model with three categories of baselines: (1) KG Reasoning Models. It consists of traditional KG reasoning models that do not consider timestamps, including DistMultyang2014embedding , ComplEx 2016Complex , ConvE 2017Convolutional , and RotatE 2019RotatE . (2) Interpolated TKG Reasoning models. We also compare our proposed model with four interpolated TKG reasoning methods, namely TTransE 2018Deriving , TA-DistMult A2018Learning , DE-SimplE 2020Diachronic , and TNTComplEx 2020Tensor . (3) Extrapolated TKG Reasoning models. We include state-of-the-art extrapolated TKG reasoning methods that infers future missing facts, including TANGO-Tucker han2021learning , TANGO-DistMult han2021learning , CyGNet zhu2021learning , RE-NET jin2020Renet , RE-GCN li2021temporal , TITer sun2021timetraveler , xERTE han2020explainable , CEN li-etal-2022-complex , GHT sun2022graph , and DaeMon dong2023adaptive .
4.1.4 Implementation Details
@@
For history temporal graph construction, we perform a grid search on the history length $m$ and present overview results with the lengths 25, 15, 10, 8, corresponding to the datasets ICEWS18, GDELT, WIKI, and YAGO in Table 3, which is described in detail presented in Figure 7. The embedding dimension $d$ is set to 64 for temporal path representation. For query-aware temporal path processing, we set the number of temporal path aggregation layers $\omega$ to 6 for ICEWS18 and GDELT datasets, and 4 for WIKI and YAGO datasets. We conduct layer normalization and shortcut on the aggregation layers. The activation function $relu$ is adopted for the aggregation of the temporal path. Similar to zhu2021neural , we use different temporal edge representations in different aggregation layers, that is the learnable parameter in $\bm{\Psi}$ and $\bm{\Upsilon}$ is independent of each aggregation iteration 10.5555/3454287.3455662 yang2017differentiable . To facilitate parameter learning, we have set the number of negative sampling to 64 and the hyperparameter $\alpha$ in the regularization term to 1. We use Adam kingma2014adam for parameter learning, with a learning rate of ${1e-4}$ for YAGO and ${5e-4}$ for others. The maximum epoch of training has been set to 20. We also design an experiment in an inductive setting, which predicts facts with unseen entities in the training set, to demonstrate the inductive ability of our proposed model. All experiments were conducted with EPYC 7742 CPU and 8 TESLA A100 GPUs.
4.2 Experimental Results (RQ1)
@@
Table 3: Overall Performance Comparison of Different Methods. Evaluation metrics are time-aware filtered MRR and Hits@{1,3,10}. All results are multiplied by 100. The best results are highlighted in bold. And the second-best results are highlighted in underline. (Higher values indicate better performance.)
@@ 4.3 Ablation Study (RQ2)
4.3.1 Effectiveness of Temporal Relation Encoding
@@
Figure 4: Ablation Results on Temporal Encoding.
@@
Table 4: Impact of the Independence Setting of Parameters in Temporal Encoder.
@@
4.3.2 Variants of Temporal Edge Merging Method
The temporal edge merging operation $\textsc{Tmsg}(\cdot)$ utilizes the message passing mechanism to iteratively extend the path length while merging the query-relevant path information. As mentioned in Equation 6, we employ the scaling operator from DistMult yang2014embedding for computing the current temporal path representation with the features of temporal edges through element-wise multiplication. Additionally, we have also conducted experiments using the translation operator from TransE bordes2013translating (i.e., element-wise summation) and the rotation operator from RotatE 2019RotatE to provide further evidence supporting the effectiveness of the temporal edge merging operation.
@@ 4.3.2 Variants of Temporal Edge Merging Method
Figure 5: Ablation Results on Temporal Edge Merging Method.
@@
Figure 6: Ablation Results on Path Aggregation Method.
4.3.3 Variants of Path Aggregation Method
The path aggregation operation $\textsc{Agg}(\cdot)$ is performed after merging the temporal edges at each step, and the aggregated feature serves as the temporal path representation from the query subject entity to the object entity candidates for future fact reasoning. As mentioned in Section 3.4.2, we employ the principal neighborhood aggregation (PNA) corso2020principal as the aggregator to aggregate the propagated messages. PNA aggregator considers mean, maximum, minimum, and standard deviation as aggregation features to obtain comprehensive neighbor information. To validate its effectiveness, we compare it with several individual aggregation operations: SUM, MEAN, and MAX.
4.4 Parameter Study (RQ3)
@@
4.4.1 Sampling History Length
In order to comprehensively capture the interconnection patterns between entities in the historical subgraphs, we constructed a history temporal graph to learn query-aware temporal paths between entities, enabling the prediction of future missing facts. Unlike previous methods based on subgraph sequences li2021temporal ; li-etal-2022-complex , we fuse multiple historical subgraphs for reasoning. However, it may lead to the inclusion of redundant edge information from different timestamps when the historical length is excessively long, resulting in unnecessary space consumption in the learning process. Moreover, facts that are far from the prediction target timestamp have minimal contribution to the prediction, even with the addition of relative distance-aware temporal encoding in the temporal relation encoder. Conversely, if the history length is too short, it can hinder the model’s ability to capture cross-time features between entities, thereby reducing the modeling performance for temporal patterns in history temporal graph. Hence, we conducted a discussion on the history length to find a balance. The experimental results of the history length study are shown in Figure 7.
@@ 4.4.1 Sampling History Length
Figure 7: The Performance of Different History Length Settings.
Table 5: Density and Interval of Datasets ($|\mathcal{V}_{avg}|$ is the average number of entities involved in the subgraph of each timestamp.).
@@
4.4.2 Number of Temporal Path Aggregation Layers
The temporal path aggregation layer is a fundamental computational unit in TiPNN. It operates on the constructed history temporal graph using the temporal edge merging and path aggregation operation to learn query-aware temporal path representations for future fact inference. The number of layers in the temporal path aggregation layer directly impacts the number of hops in the temporal message passing on the history temporal graph, which corresponds to the logical maximum distance of temporal paths. Therefore, setting the number of layers faces a similar challenge as determining the history length in Section  4.4.1.
@@ 4.4.2 Number of Temporal Path Aggregation Layers
Figure 8: The Performance of Different Numbers of Aggregation Layers.
@@
4.5 Comparison on Prediction Time (RQ4)
To further analyze the efficiency of TiPNN, we compared the inference time of TiPNN with DaeMon dong2023adaptive on the TKG reasoning task. For a fair comparison, we use the test sets of four datasets and align the model parameters under the same setting and environment. The runtime comparison between the two models is illustrated in Figure 9, where we present the ratio of their inference times in terms of multiples of the unit time. The comparative experiments demonstrate that TiPNN achieves significant reductions in runtime compared to DaeMon, with time savings of approximately 80%, 69%, 67%, and 73% on the ICEWS18, GDELT, WIKI, and YAGO datasets, respectively.
Figure 9: Runtime Comparison (The runtime is proportionally represented as multiples of the unit time.).
@@ 4.5 Comparison on Prediction Time (RQ4)
4.6 Inductive Setting (RQ5)
To validate the model’s inductive reasoning ability, we also design experiments for TiPNN in an inductive setting. The inductive setting is a common scenario in knowledge graphs, where during the training process, the model can only access a subset of entities, and during testing, it needs to reason about unseen entities Teru2020InductiveRP . This setting aims to simulate real-world situations where predictions are required for new facts based on existing knowledge. In the context of the temporal knowledge graph, the inductive setting requires the model to perform reasoning across time, meaning that the learned representations from the historical subgraphs should generalize to future timestamps for predicting missing facts. This demands the model to possess the ability to generalize to unseen entities, leveraging the learned temporal connectivity patterns from the historical subgraphs.
@@ 4.6 Inductive Setting (RQ5)
In the context of TKGs, there are few existing datasets suitable for inductive validation. Therefore, we follow the rules commonly used in KG for inductive settings and construct an inductive dataset specifically tailored for TKGs Teru2020InductiveRP . Finally, we conduct experiments on our proposed model using the inductive dataset to evaluate its inference performance under the inductive setting.
4.6.1 Inductive Datasets
@@
Table 6: Statistics of Inductive Datasets ($\mathcal{E}_{train}$, $\mathcal{E}_{valid}$, $\mathcal{E}_{test}$ are the numbers of facts in training, validation, and test sets.).
@@
4.6.2 Results of Inductive Experiment
@@
Table 7: Inductive Performance of TiPNN ($\rightarrow$ points from training set to test set corresponding to the inductive setting, and the row results without arrow marked is the corresponding transductive setting as the baseline. ‘Bias’ denotes represents the difference in comparison.).
@@
4.7 Reasoning Evidence (RQ6)
Since we have integrated historical information into the constructed history temporal graph and perform inference on future facts by modeling temporal paths, we can use the path in the history temporal graph to provide evidence for the inference process and visualize the reasoning basis. An intuitive idea is that the model should provide important reasoning paths from the history temporal graph that contribute significantly to the inference for the corresponding response. Although the temporal path representation we obtain is a comprehensive representation logically aggregated from multiple paths in the history temporal graph, we can still estimate the importance of each individual path following the local interpretation method cf736d955d6e4296a9b7255bfee3b403 ; 10.1007/978-3-319-10590-1_53 . As described in Section 3.5.1, we convert the learned temporal path representation into corresponding scores, and thus, we can estimate the importance of paths through backtracking.
Table 8: Reasoning Evidence of Responses to Selected Queries from ICEWS18. The queries, responses, and the corresponding top-2 reasoning evidence with importance (higher values indicate higher importance) are provided. The attached superscript ${}^{-1}$ denotes the inverse relations. The timestamp in the quadruple is converted into mm/dd/yyyy format for representation.
@@ 4.7 Reasoning Evidence (RQ6)
0.784 $<$Shinzo Abe, Express intent to meet or negotiate, China, 10/22/2018$>$.
@@
[…]
$<$Shinzo Abe, Make a visit, China, 10/22/2018$>$.
@@
1.613 $<$European Union, Engage in diplomatic cooperation, United Kingdom, 10/14/2018$>$.
1.044 $<$European Union, Express intent to engage in diplomatic cooperation (such as policy support),
United Kingdom, 10/22/2018$>$.
@@
[…]
@@
0.757 $<$Health Ministry (India), Make an appeal or request, Citizen (India), 10/26/2018$>$.
@@
[…]
Drawing inspiration from path interpretation zhu2022neural , we define the importance of a path as the weights assigned to it during the iteration process. We achieve this by computing the partial derivative of the temporal path scores with respect to the path’s weights. Specifically, for a reasoning response $(s,r,o,t+1)$ of a missing object query $(s,r,?,t+1)$, we consider the top-k paths as the basis for inference, which is defined as shown in Equation 16.
$\text{P}_{1},\text{P}_{2},\cdots,\text{P}_{k}=\underset{\text{P}\in\mathcal{P}_{s\rightarrow o}^{\prec t+1}}{\text{Top-}k}\frac{\partial\ p(s,r,o)}{\partial\ P}$ (16)
@@
5 Related Work
5.1 Knowledge Graph Reasoning
In recent years, knowledge graph representation learning has witnessed significant developments, aiming to embed entities and relations into continuous vector spaces while capturing their semantics bordes2013translating ; 2019RotatE ; li-etal-2022-transher . These methods can be broadly categorized into three categories: translation-based methods, semantic matching methods, and neural network-based methods. Translation-based methods, such as TransE bordes2013translating . It considers relations as translations from subject entities to object entities in the vector space. Building upon TransE, several improved methods have been proposed, including TransH wang2014knowledge , TransR lin2015learning , TransD ji2015knowledge and TransG xiao2015transg , which introduce various strategies to enhance the modeling of entity-relation interactions. For semantic matching methods, RESCAL nickel2011three proposes a tensor-based relational learning approach capable of collective learning. DistMult yang2014embedding simplifies RESCAL using diagonal matrices for efficiency, while HoIE nickel2016holographic and ComplEx trouillon2016complex extend the representation capacity by incorporating higher-order interactions and complex-valued embeddings liu2020representation , respectively. Neural network-based methods have also gained attention, with approaches like GCN welling2016semi , R-GCN schlichtkrull2017modeling , WGCN shang2019end , VR-GCN ye2019vectorized , and CompGCN vashishth2019composition , which integrate content and structural features within the graph, allowing for joint embedding of entities and relations in a relational graph Ning2021LightCAKEAL ; qiao2020context . These methods effectively capture complex patterns and structural dependencies in KGs, pushing the boundary of KG representation learning. Despite the successes of existing methods in reasoning with static KGs, they fall short when it comes to predicting temporal facts due to the lack of temporal modeling.
5.2 Temporal Knowledge Graph Reasoning
For interpolation reasoning, the objective is to infer missing facts from the past within observed data dasgupta2018hyte ; esteban2016predicting ; A2018Learning . TTransE 2018Deriving extends the TransE bordes2013translating by incorporating relations and timestamps as translation parameters between entities. TA-TransE garcia2018learning integrates the timestamps corresponding to fact occurrences into the relation representations, while HyTE dasgupta2018hyte associates timestamps with corresponding hyperplanes. Additionally, based on ComplEx trouillon2016complex , TNTComplEx 2020Tensor adopts a unique approach where the TKGs are treated as a 4th-order tensor, and the representations are learned through canonical decomposition. However, these methods are not effectively applicable for predicting future facts 2020Diachronic ; han2020dyernie ; 2018Deriving ; sadeghian2016temporal .
In contrast, the extrapolation setting, which this work focuses on, aims to predict facts in future timestamps based on historical TKG sequences. Know-Evolve trivedi2017know uses the temporal point process to represent facts in the continuous time domain. However, it falls short of capturing long-term dependencies. Similarly, DyREP trivedi2019dyrep posits representation learning as a latent mediation process and captures the changes of the observed processes. CyGNet zhu2021learning utilizes a copy-generation mechanism that collects frequently repeated events with the same subject entities and relations to the query for inferring. RE-NET jin2020Renet conducts GRU and GCN to capture the temporal and structural dependencies in a sequential manner. RE-GCN li2021temporal considers both structural dependencies and static properties of entities at the same time, modeling in evolving manner. TANGO han2021learning designs neural ordinary differential equations to represent and model TKGs for continuous-time reasoning. Most existing methods are primarily transductive in nature, only focusing on modeling the graph features based on entities while overlooking the intrinsic temporal logical rules within TKG reasoning. As an illustration, xERTE han2020explainable assumes that relations do not evolve and constructs an inference graph by sampling nodes from historical subgraphs, learning time-aware entity embeddings on it, yet neglecting complex relationship patterns. From another perspective, some path-based reasoning methods were proposed, which not only provide a new perspective for TKG reasoning but also serve as a reference for this work. For example, TITer sun2021timetraveler augments the historical subgraph sequence by adding auxiliary edges and employs reinforcement learning methods to guide path searching for inference. As a result, there is significant room for development in TKG reasoning, and a need for more efficient methods that can address the robustness and scalability requirements in real-world situations.
6 Conclusion